OpenStudy (anonymous):

evaluate the series 72+12+2+...

OpenStudy (nory):

There is a formula for the sum of an infinite geometric series, which this is. If a is the first term, and r is the common ration, then the sum to infinity is \[a/(1-r)\].

OpenStudy (nory):

What is a? And: What is r?

OpenStudy (nory):

@alexgriffin

OpenStudy (nory):

a= 72. r = 1/6. a/ (1-r) = 72/(1-5/6) = 72/(5/6) = 432/5. Do you understand how I got that?